# Efficient Composite Infrared Spectroscopy: Combining the Doubly-Harmonic Approximation with Machine Learning Potentials

### v1

@@ 1 Introduction
Building upon our previous evaluations of semiempirical extended tight-binding and machine learning approaches for vibrational spectra,20, 8, 10 this study introduces a “plug-in” composite strategy (cf. Fig. 1) designed to enhance the speed and accuracy of IR spectral predictions and provide a snapshot of current state-of-the-art’s capabilities. By combining harmonically approximated frequencies, i.e., from diagonalization of the mass-weighted Hessian, with molecular dipole moment derivatives calculated via diverse computational techniques, the goal is to achieve a balanced approach that optimizes both computational efficiency and predictive fidelity.
@@
Central to our methodological framework is the integration of a recently reported machine learning potential (MLP), MACE-OFF2321 for the computation of molecular frequencies. This model employs the MACE22 MLP architecture for organic molecules based on accurate numerical data, promising to bridge the accuracy gap typically associated with faster computational methods. The current implementation of this MLP does not include molecular electrostatics and so obtaining a dipole moment is not possible using only MACE-OFF23. Therefore, a separate MACE model is introduced and trained for dipole moment prediction. The kQEq model23, 24 was also considered. We compare MLP-derived dipole moments with those obtained from classical atomic charge models and more traditional quantum mechanically derived dipole moments, providing a comprehensive evaluation of the performance for each method. Additionally, the ability to obtain the correct molecular geometries, i.e. find the minima on the potential energy landscape via optimization procedures,25, 26 plays a crucial role in a composite approach, influencing both frequency calculations and dipole moment predictions. By assessing method performance across these method capabilities, we not only seek to validate the utility of MLPs in theoretical spectroscopy, but also explore their integration into established computational workflows. Some further key targets are to investigate whether ML methods trained on the SPICE dataset27 are suitable surrogates for predictions at DFT-level accuracy, and whether their composite combination with other, for example SQM methods can advance our predictive capabilities in general.
@@ 2.1 The double-harmonic approximation
In the so-called ‘double-harmonic approximation’, which gets its name from the Taylor series truncation after the quadratic terms for both frequencies and dipole derivatives,28, 5 the intensity $A_{\tilde{\nu_{p}}}$ for the normal mode $p$ is assumed proportional to the squared dipole derivative of the corresponding normal mode and the integral absorption coefficient is then given as
@@ 2.2 Computational details
For the computation of molecular dipole moments, molecular geometries, frequencies and IR spectra, we generally refer to a benchmark set of 7193 gas-phase molecules20 ranging from three to 77 atoms in size, and containing the elements H, C, N, O, F, Cl, Br, S, and P, which coincides with the elements available in MACE-OFF23.21 For simplicity, we refer to this set as ‘IR7193’ in the following. The corresponding molecules can generally be regarded as ‘drug-like‘ with regards to the average size and chemical composition. This set was originally composed20 simply by obtaining all freely available experimental gas-phase reference IR spectra from the NIST database29 which were then compared with computed spectra for the semiempirical GFNn-xTB approach,12 the force-field GFN-FF,30 and DFT B3LYP-3c (an abbreviation for B3LYP-D3(BJ)ATM-gCP/def2-mSVP31, 32, 33, 34, 35). B3LYP-3c herein was identified as a well-balanced DFT level of theory for IR spectra computation and can serve as a theoretical reference for testing other low-cost methods. To avoid additional conformational effects,20 and effects of possible experimental measurement impurities, we take the readily available theoretical gas-phase spectra calculated at B3LYP-3c level as our reference.
@@
Analysis tools for the combination of computed Hessian matrices and dipole derivatives into IR spectra at https://github.com/pprcht/IRtools, as well as for the comparison of IR spectra can be found at https://github.com/grimme-lab/newspecmatch. Our comparison of IR spectra is based on a similarity ‘match’score $r_{\mathrm{msc}}$ based on the Cauchy-Schwarz inequality as defined by Eq. S3 in the supporting information and can have values from 0 (no match) to 1 (perfect correlation). For more information on spectra comparison with this metric and alternative measures (Euclidean norm $r_{\mathrm{euc}}$, the Spearman correlation coefficient $r_{\mathrm{scc}}$, Pearson correlation coefficient $r_{\mathrm{pcc}}$) we refer the reader to the literature,42, 43, 44, 45, 20, 3, 46 and the supplementary material.
@@ 3.1 Performance for geometry optimization
The performance of a given level of theory for geometry optimization is important, because frequencies are obtained in the harmonic approximation, where the molecule must correspond to a minimum on the energy landscape. If the molecular structure is not a minimum, or converges to a different conformation, identified by a high root-mean-square-deviation (RMSD) of atomic coordinates,38 imaginary frequencies or incorrect fundamental modes will be observed, respectively. Hence, having a well-defined geometry reference is necessary. It is generally recognized that structure optimizations at the DFT level are much less sensitive to the choice of functional than, for example, energy and property calculations.47 Commonly, mGGA functionals are sufficient for geometry optimisation, where the general recommendation is to use basis sets of triple-$\zeta$ quality, or smaller basis sets in combination with corrections for the basis set superposition error (BSSE).34, 47 Reference geometries for the IR7193 set are available at the B3LYP-3c level of theory which, as a hybrid functional with BSSE-corrected basis set, is more than sufficient to compare force-field, SQM, and MLP results. Geometry optimizations for all low-cost methods were herein started from the B3LYP-3c optimized structure in order to ensure convergence to the closest minima on the respective potential energy landscape.
The GFN1-xTB and GFN2-xTB levels of theory along with the GFN-FF and MACE-OFF23 force fields were used to optimize and compare structures in the database. Tables 1 and 2 summarise the corresponding RMSD statistics with regard to the B3LYP-3c reference structures. Figure 2 illustrates the distribution of RMSDs as histograms compared to a log-normal distribution.
@@ 3.2 Performance for vibrational frequencies
Having discussed performance for optimized geometries in the previous section, we now turn our attention to the calculation of properties defining the IR spectra prediction, with a primary focus on the calculation of internal molecular vibrations in the harmonic approximation. A simple assessment of frequency quality is possible via comparison of the harmonic zero-point vibrational energy (ZPVE), which is approximated48 as half the sum over all frequencies
@@
The ZPVE together with enthalpic and entropic contributions, produces the thermodynamic contributions to the Gibbs free energy in supramolecular calculations.49 Although Eq. 7 is not precise due to neglect of anharmonicity48 and a significant amount of information is condensed into the single value of the ZPVE, it offers the distinct advantage of being independent of the permutational order of fundamental frequencies, which can vary with the level of theory used and complicate IR spectra comparison. Furthermore, it allows for the direct derivation of frequency scaling factors, which are commonly applied in (harmonic) DFT frequency calculations to counteract the effects of inadequate descriptions of vibrational anharmonicity.50, 48, 51, 52, 53
The ZPVE, calculated for all molecules in the IR7193 set at the GFN1- and GFN2-xTB, the GFN-FF, and the MACE-OFF23 levels of theory, was evaluated against reference values obtained from B3LYP-3c. To extend the comparison, modified frequencies from a scaled version of B3LYP-3c were used to obtain linear frequency scaling factors, by taking the corresponding fraction of ZPVEs and averaging over the whole test set for each method. The modified B3LYP-3c results, denoted as B3LYP-3c(mscal), employ a special non-linear scaling scheme following Ref. 20 and the work of Pulay et al.54, 55 The corresponding “mass-scaling” factors were fitted to the experimental IR spectra and should account for a large part of the vibrational anharmonicity.56, 20 Using the B3LYP-3c(mscal) as a reference, the ZPVEs were evaluated twice: Once, using the raw ZPVE obtained with each low-cost potential, and once with the ZPVE of each low-cost method based on frequencies scaled by the linear factor determined from the aforementioned fraction. Error measures (mean deviation, mean absolute error, root-mean-square error, and standard deviation) for the investigated levels of theory are shown in Table 3.
Table 3: Comparison low-cost potentials for calculation of ZPVE. Three sets of data are shown: First values with plain B3LYP-3c as a reference, second B3LYP-3c mass-scaled frequencies as a reference (cf. Ref. 20), and finally results with linear scaling factors applied and mass-scaled B3LYP-3c as a reference. Values are given in kcal  mol-1.
@@
The final block of data in Tab. 3 provides insights into the linear scaling factors determined using the mass-scaled B3LYP-3c ZPVE as a reference. Notably, the MACE-OFF23 models have consistent scaling factors around 0.96, which is close to the literature scaling factor of 0.97 for B3LYP.48, 52, 53 MAEs for the MACE MLP are lowered to just 0.22 to 0.23 kcal  mol-1, while RMSEs are lowered to 0.30 to 0.31 kcal  mol-1. While, to the best of our knowledge, no scaling factors have been investigated for the SPICE $\omega$B97M-D3 reference, the related $\omega$B97X-D3 functional has known scaling factors for fundamental frequencies of around 0.95,53 which is close to the observed MACE-OFF23 scaling factors. This result suggests that the MLP behaves similarly to a hybrid DFT method, supporting the previous assessment of its performance. Importantly, a significant harmonic frequency scaling factor shows a systematic error in particular for high-frequency vibrational modes as these dominate the harmonic ZPVE. On the other hand, the GFN methods, being semiempirical and force-field methods, do not exhibit systematic scaling factors, reflecting the less consistent behavior in scaling vibrational frequencies that is known for these levels of theory in the literature.57, 58, 3 Nevertheless, the GFN methods MAE is lowered slightly to 1.01 to 1.118 kcal  mol-1, and the respective RMSE is lowered to 1.34 to 1.62 kcal  mol-1. Overall, the consistency in the MACE models underscores their reliability and further validates their use for molecular frequency calculations within the harmonic approximation, aligning closely with hybrid DFT methodologies. Nonetheless, the GFN methods may have an advantage in computational efficiency for drug sized molecules, as addressed in the following.
@@ 3.3 Performance for molecular dipole moments
IR intensities are related to the molecular dipole moment $\mu$, as described by Eq. 5. Therefore, if attempting to switch out the method used for the calculation of $\partial\mu/\partial Q_{p}$ in a composite approach, it is sensible to benchmark the performance for $\mu$. Employing again the IR7193 set of molecules in addition to three conformational benchmark sets (MALT222,59 MPCONF196,60 and 37conf861), this comparison was done for the B3LYP-3c level of DFT, the semiempirical GFN1- and GFN2-xTB Hamiltonians, the CEH method,40 an empirical EEQ model using parameter sets from the D4 dispersion correction and the GFN-FF force-field, as well as a novel MLP for dipole prediction, MACE-$\mu$, and a new fit of the kQEq model23, 24 for molecules containing elements HCNOF. Both MACE-$\mu$ and kQEq were trained on the SPICE dataset27 for this study. The $\omega$B97M-D3(BJ)/def2-TZVPPD level of theory was used as a reference for these dipole calculations, and is also used as the reference in the SPICE dataset. The $\omega$B97M-D3(BJ) level itself was benchmarked for dipole prediction on a much smaller set of molecules, originally composed by Head-Gordon et al.62, 63 and refers to CCSD(T)/CBS data, which can be found in the supporting information.
Since IR intensities are calculated from derivatives of the dipole moment (Eq. 6) rather than the dipole moment directly, two criteria are important for the prediction of $\mu$: first, the total dipole moment, and second the orientation of the dipole vector. The total dipole moment is straightforward to benchmark, where we employ the regularized dipole metric $\Delta\mu={|\mu-\mu_{\mathrm{ref}}|}/{\max{\{\mu_{\mathrm{ref}}},1\leavevmode\nobreak\ \mathrm{Debye}\}}$ proposed by Hait and Head-Gordon,62 in order to handle small and large dipole moments on a more even footing. The corresponding results are shown as violin plots for the summarized IR7193, MALT222, MPCONF196 and 37conf8 sets in Fig. 5 and Table 4. Additional information is provided in the supporting information.
Figure 5: Violin plots for regularized errors of dipole moments at a given level of theory with regard to the reference $\omega$B97M-D3(BJ) dipole moments. Note, kQEq was evaluated for the subset containing HCNOF elements only.
@@
Table 4: Errors for regularized dipole moments at a given level of theory. Reference dipole moments were obtained at the $\omega$B97M-D3(BJ) level. Note, kQEq was evaluated for the subset containing HCNOF elements only. All values are dimensionless.
@@
Several observations can be made for the dipole moment calculations. B3LYP-3c, which serves as our reference for IR spectra, molecular geometries and ZPVE, shows performance very close to the $\omega$B97M-D3(BJ) reference with a regularized RMSE of just 6.97%. Furthermore, as can be seen from Fig. 6, B3LYP-3c and the high-level DFT reference also exhibit the best dipole vector alignment with a mean close to unity and a standard deviation of just 0.08. Dipole moments at this level of theory on average seem underestimated, leading to the MD of -3.25%. A probable origin of this error is the small def2-mSVP basis set employed here.63
@@
GFN1-xTB and the CEH model dipole moments are based on the distribution of Mulliken atomic charges.36, 40 Unsurprisingly, both methods exhibit similar performance for total dipole moments, with a regularized RMSE of 0.2449 and 0.2488, respectively. Characteristically, GFN1-xTB has two maxima in the violin probability density of Figure 5, resulting in an up-shifted mean compared to CEH and a significant number of molecules have strongly overestimated dipole moments. Both methods contain several outliers, including values overestimated by more than a factor of two according to the regularized dipole moments. The alignment of dipole vectors of both methods with the reference is likewise similar, although significantly worse than B3LYP-3c. Overall, GFN1-xTB slightly outperforms CEH. This result can likely be attributed to the SCC procedure in GFN1-xTB allowing a better adaptation of charge density, while CEH is a “single-shot” diagonalization (Hückel) method, lacking self-consistency.
@@
GFN2-xTB improves upon both Mulliken charge based models by atom centered dipole and multipole moments.37 While this description leads to a narrower probability distribution and better mean in Figure 5 with a RMSE of 0.1877, dipole alignment is on average worse than with either GFN1-xTB or CEH. Taking this result into account, all SQM methods offer similar robust and efficient dipole moment calculations.
@@
MACE-$\mu$, a MACE model trained on the SPICE27 dataset to predict the molecular dipole vector for this work, shows some interesting features. The model presented here formally refers to a “medium” MACE model size, some additional data for a “small” model can be found in the Supporting Information. The MD of 0.0128, the MAD of 0.0565 and RMSD of 0.0931, as well as the dipole vector alignment (cf. Fig. 6), show a generally better performance of MACE-$\mu$ than those of the SQM-based methods. Instead, the MACE-$\mu$ performance much more closely resembles that of a DFT method, although some greater outliers can be seen in the violin plots of Figure 5. Notably, our dipole moment reference method is the same as the level of theory employed in the SPICE dataset and, in fact, a number of predictions match the $\omega$B97M-D3 reference precisely. At the present time it is unclear what causes the remaining dipole moment outliers of MACE-$\mu$, but still we see it as a slightly better (although for small systems more expensive) alternative to the GFNn-xTB dipole predictions.
@@
The worst molecular dipole moments come from the classical EEQ models. The RMSE, 0.8812 and 0.8351 for the D4 and GFN-FF parameter set, respectively, is more than an order of magnitude worse than for B3LYP-3c, and roughly a factor of four worse than with the SQM methods. This is in line with previous observations for dipole moments calculated from EEQ models in the literature.23, 20 The performance can apparently be influenced by the parametrization, and the GFN-FF EEQ model, on average, shows marginally better results than the D4 EEQ parameter set. Unsurprisingly, the largest outlier, with a dipole moment overestimated by roughly a factor 10, is obtained with EEQ(D4). For the dipole vector orientation both EEQ parameter sets perform equally badly, giving broad distributions in Figure 6 that extend even into the negative regime of $\hat{\mu}\cdot\hat{\mu}^{\mathrm{ref}}$. Hence, a non-negligible number of molecules have dipole moments pointing into the wrong direction at this level of theory, two of which are shown in Figures 7a and 7b. In these cases, EEQ models clearly provide an insufficient description of the charge distribution, which can only be recovered with electronic structure calculations or by certain MLPs.23 EEQ models are prone to unphysical charge-transfer problems,64, 23, 12 which contributes to these issues.
@@
Interestingly, the kQEq model (which is an ML augmented EEQ model) performs significantly better than its classical counterparts. In terms of the dipole moments it approaches the accuracy of GFN1-xTB, although several significant outliers lead to a somewhat larger RMSE. The dipole orientations are also improved relative to the classical models, although a few outliers with dipoles pointing in the wrong direction remain. Overall, environment dependent electronegativities as used in kQEq thus improve the performance of EEQ-based models, but the underlying limitations (i.e. using a point-charge representation of the density and unphysical charge transfer effects of EEQ) are not fully cured.
@@
Figure 7c shows the aforementioned high EEQ dipole moment outlier. A similar linear molecule has been discussed, for example, in Ref. 23. Interestingly, this outlier cannot be attributed to the partial charge distribution of the carbon atoms, as illustrated by Figure S3 in the supporting information. In fact, all atomic charges at the EEQ(GFN-FF) level are very similar to those at the GFN2-xTB level, despite the latter calculation exhibiting a qualitatively well-defined dipole moment that differs from the high-level DFT reference by just 0.204 D (or 13%, referring to the regularized dipole measure). On the other hand, partial charges at the EEQ(D4) level are different from both GFN2-xTB and EEQ(GFN-FF), as well as the DFT references. A clearly wrong assignment of charges that would be reflected in the dipole moment prediction cannot be discerned in either case, although partial charges calculated at the high quality reference $\omega$B97M-D3(BJ) differ distinctively from all other methods. The severely overestimated dipole moments via the EEQ models can therefore not be unequivocally attributed to charge transfer issues from polarization or electron delocalization. A delicate balance between different contributions to the molecular dipole moment is needed for useful performance.

### v2

@@ 1 Introduction
Building upon our previous evaluations of semiempirical extended tight-binding and machine learning approaches for vibrational spectra,21, 9, 11 this study introduces a “plug-in” composite strategy (cf. Fig. 1) designed to enhance the speed and accuracy of IR spectral predictions and provide a snapshot of current state-of-the-art’s capabilities. By combining harmonically approximated frequencies, i.e., from diagonalization of the mass-weighted Hessian, with molecular dipole moment derivatives calculated via diverse computational techniques, the goal is to achieve a balanced approach that optimizes both computational efficiency and predictive fidelity.
@@
Central to our methodological framework is the integration of a recently reported MLP, MACE-OFF2322 for the computation of molecular frequencies. This model employs the MACE23 MLP architecture for organic molecules based on accurate numerical data, promising to bridge the accuracy gap typically associated with faster computational methods. The current implementation of this MLP does not include molecular electrostatics and so obtaining a dipole moment is not possible using only MACE-OFF23. Therefore, a separate MACE model is introduced and trained for dipole moment prediction. The kQEq model24, 25 was also considered. We compare MLP-derived dipole moments with those obtained from classical atomic charge models and more traditional quantum mechanically derived dipole moments, providing a comprehensive evaluation of the performance for each method. Additionally, the ability to obtain the correct molecular geometries, i.e. find the minima on the potential energy landscape via optimization procedures,26, 27 plays a crucial role in a composite approach, influencing both frequency calculations and dipole moment predictions. By assessing method performance across these method capabilities, we not only seek to validate the utility of MLPs in theoretical spectroscopy, but also explore their integration into established computational workflows. Some further key targets are to investigate whether ML methods trained on the SPICE dataset28 are suitable surrogates for predictions at DFT-level accuracy, and whether their composite combination with other, for example SQM methods can advance our predictive capabilities in general.
@@ 2.1 The double-harmonic approximation
In the so-called ‘double-harmonic approximation’, which gets its name from the Taylor series truncation after the quadratic terms for both frequencies and dipole derivatives,29, 5 the intensity $A_{\tilde{\nu_{p}}}$ for the normal mode $p$ is assumed proportional to the squared dipole derivative of the corresponding normal mode and the integral absorption coefficient is then given as
@@ 2.2 Computational details
For the computation of molecular dipole moments, molecular geometries, frequencies and IR spectra, we generally refer to a benchmark set of 7193 gas-phase molecules21 ranging from three to 77 atoms in size, and containing the elements H, C, N, O, F, Cl, Br, S, and P, which coincides with the elements available in MACE-OFF23.22 For simplicity, we refer to this set as ‘IR7193’ in the following. The corresponding molecules can generally be regarded as ‘drug-like‘ with regards to the average size and chemical composition. This set was originally composed21 simply by obtaining all freely available experimental gas-phase reference IR spectra from the NIST database30 which were then compared with computed spectra for the semiempirical GFNn-xTB approach,13 the force-field GFN-FF,31 and DFT B3LYP-3c (an abbreviation for B3LYP-D3(BJ)ATM-gCP/def2-mSVP32, 33, 34, 35, 36). B3LYP-3c herein was identified as a well-balanced DFT level of theory for IR spectra computation and can serve as a theoretical reference for testing other low-cost methods. To avoid additional conformational effects,21 and effects of possible experimental measurement impurities, we take the readily available theoretical gas-phase spectra calculated at B3LYP-3c level as our reference.
@@
Analysis tools for the combination of computed Hessian matrices and dipole derivatives into IR spectra can be found at https://github.com/pprcht/vibtools. Our comparison of IR spectra is based on a similarity ‘match’score $r_{\mathrm{msc}}$ based on the Cauchy-Schwarz inequality as defined by Eq. S3 in the supporting information and can have values from 0 (no match) to 1 (perfect correlation). For more information on spectra comparison with this metric and alternative measures (Euclidean norm $r_{\mathrm{euc}}$, the Spearman correlation coefficient $r_{\mathrm{scc}}$, Pearson correlation coefficient $r_{\mathrm{pcc}}$) we refer the reader to the literature,43, 44, 45, 46, 21, 3, 47 and the supplementary material.
@@ 3.1 Performance for geometry optimization
The performance of a given level of theory for geometry optimization is important, because frequencies are obtained in the harmonic approximation, where the molecule must correspond to a minimum on the energy landscape. If the molecular structure is not a minimum, or converges to a different conformation, identified by a high root-mean-square-deviation (RMSD) of atomic coordinates,39 imaginary frequencies or incorrect fundamental modes will be observed, respectively. Hence, having a well-defined geometry reference is necessary. It is generally recognized that structure optimizations at the DFT level are much less sensitive to the choice of functional than, for example, energy and property calculations.48 Commonly, mGGA functionals are sufficient for geometry optimisation, where the general recommendation is to use basis sets of triple-$\zeta$ quality, or smaller basis sets in combination with corrections for the basis set superposition error (BSSE).35, 48 Reference geometries for the IR7193 set are available at the B3LYP-3c level of theory which, as a hybrid functional with BSSE-corrected basis set, is more than sufficient to compare force-field, SQM, and MLP results. Geometry optimizations for all low-cost methods were herein started from the B3LYP-3c optimized structure in order to ensure convergence to the closest minima on the respective potential energy landscape.
The GFN1-xTB and GFN2-xTB levels of theory along with the GFN-FF and MACE-OFF23 force-fields were used to optimize and compare structures in the database. Tables 1 and 2 summarise the corresponding RMSD statistics with regard to the B3LYP-3c reference structures. Figure 2 illustrates the distribution of RMSDs as histograms compared to a log-normal distribution.
@@ 3.2 Performance for vibrational frequencies
Having discussed performance for optimized geometries in the previous section, we now turn our attention to the calculation of properties defining the IR spectra prediction, with a primary focus on the calculation of internal molecular vibrations in the harmonic approximation. A simple assessment of frequency quality is possible via comparison of the harmonic zero-point vibrational energy (ZPVE), which is approximated49 as half the sum over all frequencies
@@
The ZPVE together with enthalpic and entropic contributions, produces the thermodynamic contributions to the Gibbs free energy in supramolecular calculations.50 Although Eq. 7 is not precise due to neglect of anharmonicity49 and a significant amount of information is condensed into the single value of the ZPVE, it offers the distinct advantage of being independent of the permutational order of fundamental frequencies, which can vary with the level of theory used and complicate IR spectra comparison. Furthermore, it allows for the direct derivation of frequency scaling factors, which are commonly applied in (harmonic) DFT frequency calculations to counteract the effects of inadequate descriptions of vibrational anharmonicity.51, 49, 52, 53, 54
The ZPVE, calculated for all molecules in the IR7193 set at the GFN1- and GFN2-xTB, the GFN-FF, and the MACE-OFF23 levels of theory, was evaluated against reference values obtained from B3LYP-3c. To extend the comparison, modified frequencies from a scaled version of B3LYP-3c were used to obtain linear frequency scaling factors, by taking the corresponding fraction of ZPVEs and averaging over the whole test set for each method. The modified B3LYP-3c results, denoted as B3LYP-3c(mscal), employ a special non-linear scaling scheme following Ref. 21 and the work of Pulay et al.55, 56 The corresponding “mass-scaling” factors were fitted to the experimental IR spectra and should account for a large part of the vibrational anharmonicity.57, 21 Using the B3LYP-3c(mscal) as a reference, the ZPVEs were evaluated twice: Once, using the raw ZPVE obtained with each low-cost potential, and once with the ZPVE of each low-cost method based on frequencies scaled by the linear factor determined from the aforementioned fraction. Error measures (mean deviation, mean absolute error, root-mean-square error, and standard deviation) for the investigated levels of theory are shown in Table 3.
Table 3: Comparison low-cost potentials for calculation of ZPVE. Three sets of data are shown: First values with plain B3LYP-3c as a reference, second B3LYP-3c mass-scaled frequencies as a reference (cf. Ref. 21), and finally results with linear scaling factors applied and mass-scaled B3LYP-3c as a reference. Values are given in kcal  mol-1.
@@
The final block of data in Tab. 3 provides insights into the linear scaling factors determined using the mass-scaled B3LYP-3c ZPVE as a reference. Notably, the MACE-OFF23 models have consistent scaling factors around 0.96, which is close to the literature scaling factor of 0.97 for B3LYP.49, 53, 54 MAEs for the MACE MLP are lowered to just 0.22 to 0.23 kcal  mol-1, while RMSEs are lowered to 0.30 to 0.31 kcal  mol-1. While, to the best of our knowledge, no scaling factors have been investigated for the SPICE $\omega$B97M-D3 reference, the related $\omega$B97X-D3 functional has known scaling factors for fundamental frequencies of around 0.95,54 which is close to the observed MACE-OFF23 scaling factors. This result suggests that the MLP behaves similarly to a hybrid DFT method, supporting the previous assessment of its performance. Importantly, a significant harmonic frequency scaling factor shows a systematic error in particular for high-frequency vibrational modes as these dominate the harmonic ZPVE. On the other hand, the GFN methods, being semiempirical and force-field methods, do not exhibit systematic scaling factors, reflecting the less consistent behavior in scaling vibrational frequencies that is known for these levels of theory in the literature.58, 59, 3 Nevertheless, the GFN methods MAE is lowered slightly to 1.01 to 1.118 kcal  mol-1, and the respective RMSE is lowered to 1.34 to 1.62 kcal  mol-1. Overall, the consistency in the MACE models underscores their reliability and further validates their use for molecular frequency calculations within the harmonic approximation, aligning closely with hybrid DFT methodologies. Nonetheless, the GFN methods may have an advantage in computational efficiency for drug sized molecules, as addressed in the following.
@@ 3.3 Performance for molecular dipole moments
IR intensities are related to the molecular dipole moment $\mu$, as described by Eq. 5. Therefore, if attempting to switch out the method used for the calculation of $\partial\mu/\partial Q_{p}$ in a composite approach, it is sensible to benchmark the performance for $\mu$. Employing again the IR7193 set of molecules in addition to three conformational benchmark sets (MALT222,60 MPCONF196,61 and 37conf862), this comparison was done for the B3LYP-3c level of DFT, the semiempirical GFN1- and GFN2-xTB Hamiltonians, the CEH method,41 an empirical EEQ model using parameter sets from the D4 dispersion correction and the GFN-FF force-field, as well as a novel MLP for dipole prediction, MACE-$\mu$, and a new fit of the kQEq model24, 25 for molecules containing the elements HCNOF. Both MACE-$\mu$ and kQEq were trained on the SPICE dataset28 for this study. The $\omega$B97M-D3(BJ)/def2-TZVPPD level of theory was used as a reference for these dipole calculations, and is also used as the reference in the SPICE dataset. The $\omega$B97M-D3(BJ) level itself was benchmarked for dipole prediction on a much smaller set of molecules, originally composed by Head-Gordon et al.63, 64 and refers to CCSD(T)/CBS data, which can be found in the supporting information.
Since IR intensities are calculated from derivatives of the dipole moment (Eq. 6) rather than the dipole moment directly, two criteria are important for the prediction of $\mu$: first, the total dipole moment, and second the orientation of the dipole vector. The total dipole moment is straightforward to benchmark, where we employ the regularized dipole metric $\Delta\mu={|\mu-\mu_{\mathrm{ref}}|}/{\max{\{\mu_{\mathrm{ref}}},1\leavevmode\nobreak\ \mathrm{Debye}\}}$ proposed by Hait and Head-Gordon,63 in order to handle small and large dipole moments on a more even footing. The corresponding results are shown as violin plots for the summarized IR7193, MALT222, MPCONF196 and 37conf8 sets in Fig. 5 and Table 4. Additional information is provided in the supporting information.
Figure 5: Violin plots for regularized errors of dipole moments at a given level of theory with regard to the reference $\omega$B97M-D3(BJ) dipole moments. Note, kQEq was evaluated for the subset of molecules containing HCNOF elements only.
@@
Table 4: Errors for regularized dipole moments at a given level of theory. Reference dipole moments were obtained at the $\omega$B97M-D3(BJ) level. Note, kQEq was evaluated for the subset of molecules containing HCNOF elements only. All values are dimensionless.
@@
Several observations can be made for the dipole moment calculations. B3LYP-3c, which serves as our reference for IR spectra, molecular geometries and ZPVE, shows performance very close to the $\omega$B97M-D3(BJ) reference with a regularized RMSE of just 6.97%. Furthermore, as can be seen from Fig. 6, B3LYP-3c and the high-level DFT reference also exhibit the best dipole vector alignment with a mean close to unity and a standard deviation of just 0.08. Dipole moments at this level of theory on average seem underestimated, leading to the MD of -3.25%. A probable origin of this error is the small def2-mSVP basis set employed here.64
@@
GFN1-xTB and the CEH model dipole moments are based on the distribution of Mulliken atomic charges.37, 41 Unsurprisingly, both methods exhibit similar performance for total dipole moments, with a regularized RMSE of 0.2449 and 0.2488, respectively. Characteristically, GFN1-xTB has two maxima in the violin probability density of Figure 5, resulting in an up-shifted mean compared to CEH and a significant number of molecules have strongly overestimated dipole moments. Both methods contain several outliers, including values overestimated by more than a factor of two according to the regularized dipole moments. The alignment of dipole vectors of both methods with the reference is likewise similar, although significantly worse than B3LYP-3c. Overall, GFN1-xTB slightly outperforms CEH. This result can likely be attributed to the SCC procedure in GFN1-xTB allowing a better adaptation of charge density, while CEH is a “single-shot” diagonalization (Hückel) method, lacking self-consistency.
@@
GFN2-xTB improves upon both Mulliken charge based models by atom centered dipole and multipole moments.38 While this description leads to a narrower probability distribution and better mean in Figure 5 with a RMSE of 0.1877, dipole alignment is on average worse than with either GFN1-xTB or CEH. Taking this result into account, all SQM methods offer similar robust and efficient dipole moment calculations.
@@
MACE-$\mu$, a MACE model trained on the SPICE28 dataset to predict the molecular dipole vector for this work, shows some interesting features. The model presented here formally refers to a “medium” MACE model size, some additional data for a “small” model can be found in the Supporting Information. The MD of 0.0128, the MAD of 0.0565 and RMSD of 0.0931, as well as the dipole vector alignment (cf. Fig. 6), show a generally better performance of MACE-$\mu$ than those of the SQM-based methods. Instead, the MACE-$\mu$ performance much more closely resembles that of a DFT method, although some greater outliers can be seen in the violin plots of Figure 5. Notably, our dipole moment reference method is the same as the level of theory employed in the SPICE dataset and, in fact, a number of predictions match the $\omega$B97M-D3 reference precisely. At the present time it is unclear what causes the remaining dipole moment outliers of MACE-$\mu$, but still we see it as a slightly better (although for small systems more expensive) alternative to the GFNn-xTB dipole predictions.
@@
The worst molecular dipole moments come from the classical EEQ models. The RMSE, 0.8812 and 0.8351 for the D4 and GFN-FF parameter set, respectively, is more than an order of magnitude worse than for B3LYP-3c, and roughly a factor of four worse than with the SQM methods. This is in line with previous observations for dipole moments calculated from EEQ models in the literature.24, 21 The performance can apparently be influenced by the parametrization, and the GFN-FF EEQ model, on average, shows marginally better results than the D4 EEQ parameter set. Unsurprisingly, the largest outlier, with a dipole moment overestimated by roughly a factor 10, is obtained with EEQ(D4). For the dipole vector orientation both EEQ parameter sets perform equally badly, giving broad distributions in Figure 6 that extend even into the negative regime of $\hat{\mu}\cdot\hat{\mu}^{\mathrm{ref}}$. Hence, a non-negligible number of molecules have dipole moments pointing into the wrong direction at this level of theory, two of which are shown in Figures 7a and 7b. In these cases, EEQ models clearly provide an insufficient description of the charge distribution, which can only be recovered with electronic structure calculations or by certain MLPs.24 EEQ models are prone to unphysical charge-transfer problems,65, 24, 13 which contributes to these issues.
@@
Interestingly, the kQEq model (which is an ML augmented EEQ model) performs significantly better than its classical counterparts. In terms of the dipole moments it approaches the accuracy of GFN1-xTB, although several significant outliers lead to a somewhat larger RMSE. The dipole orientations are also improved relative to the classical models, although a few outliers with dipoles pointing in the wrong direction remain. Overall, environment dependent electronegativities, as used in kQEq, improve the performance of EEQ-based models, but the underlying limitations (i.e. using a point-charge representation of the density and unphysical charge transfer effects of EEQ) are not fully cured.
@@
Figure 7c shows the aforementioned high EEQ dipole moment outlier. A similar linear molecule has been discussed, for example, in Ref. 24. Interestingly, this outlier cannot be attributed to the partial charge distribution of the carbon atoms, as illustrated by Figure S3 in the supporting information. In fact, all atomic charges at the EEQ(GFN-FF) level are very similar to those at the GFN2-xTB level, despite the latter calculation exhibiting a qualitatively well-defined dipole moment that differs from the high-level DFT reference by just 0.204 D (or 13%, referring to the regularized dipole measure). On the other hand, partial charges at the EEQ(D4) level are different from both GFN2-xTB and EEQ(GFN-FF), as well as the DFT references. A clearly wrong assignment of charges that would be reflected in the dipole moment prediction cannot be discerned in either case, although partial charges calculated at the high quality reference $\omega$B97M-D3(BJ) differ distinctively from all other methods. The severely overestimated dipole moments via the EEQ models can therefore not be unequivocally attributed to charge transfer issues from polarization or electron delocalization. A delicate balance between different contributions to the molecular dipole moment is needed for useful performance.